could you please help me on this problem.
the difference between two complimentary angles is 32. find the larger angle.
ive started
c-2=32
don't know if equation is right

Answer by bluemockingjay7(41)

This is the correct solution:
smaller = x
larger = 90 - x
90 - x - x = 32
90 - 2x = 32
-2x = 32 - 90
-2x = -58
-2x/-2 = -58/-2
x = 29
smaller angle is 29 degrees
larger angle is 90 - 29 = 61 degrees
check:
61 + 29 = 90 degrees = complementary
61 - 29 = 32 degrees = difference